Dave Bowers collects U.S. gold coins. He has a collection of 46 coins. Some are $10 coins and the rest are $20 coins. If the face value of the coin is $640, then how many of each denoination does he have? Use the six-step method.
Answer by venugopalramana(3286)

Dave Bowers collects U.S. gold coins. He has a collection of 46 coins. Some are $10 coins( X SAY ) and the rest are $20 coins ( Y SAY ).
X+Y=46..............OR...........Y=46-X..................1
If the face value of the coin is $640,
F.V.OF $10 COINS = 10X
F.V.OF $20 COINS = 20Y
TOTAL VALUE = 10X+20Y=10X+20(46-X)=640
10X+920-20X=640
10X=920-640=280
X=28
Y=46-28=18
HENCE $ 10 COINS ARE 28 IN NUMBER AND
$ 20 COINS ARE 18 IN NUMBER



then how many of each denoination does he have? Use the six-step method.